Multi-part question




Your answer is incorrect. Try again. Determine the overhead allocation for the family model and the deluxe model using three activity cost pools. Oerhead Allocation Family model Deluxe model What is the difference in allocation between two activity cost pools and three activity cost pools? Is the difference in allocation worth using the third activity cost pool? Difference The additional cost pool the overhead cost allocation significantly. Click if you would like to Show Work for this question: Open Show Work
Answer #1

1.Using traditional approach, overhead rate is based on direct labor cost

= Total overheads/Total direct labor cost

= 468,000/234,000 = 200%

Or $2 per dollar of direct labor cost

2.

Machining = 280,800/4,000

= $70.2 per hour

Machine Set up = 187,200/1,000

= $187.2 per set up

Overhead Amount allocated under ABC

Family Model = 2,000*70.2 + 200*187.2 = $177,840

Deluxe Model = 2,000*70.2 + 800*187.2 = $290,160

Difference

Family

Deluxe

Traditional

156,000

312,000

ABC

177,840

290,160

Difference

(21,840)

21,840

Machining = 249,600/4,000 = $62.4 per hour

Set up = 176,800/1000 = $176.8 per setup

Packaging = 41,600/125 = $332.8 per hour

Allocation:

family Model = 2,000*62.4 + 200*176.8 + 50*332.8 = $176,800

Deluxe Model = 2,000*62.4 + 800*176.8 + 75*332.8 = $291,200

Difference in allocation = $1,040

The additional cost pool DOES NOT change the overhead cost allocation significantly